OpenStudy (anonymous):

i need help with the file i add pl look at it help me I WILL AWARD MEDAL!!!!

OpenStudy (anonymous):

i need help with the file i add pl look at it help me I WILL AWARD MEDAL!!!!

OpenStudy (anonymous):

a) m=$0.28,b=$163.90 b)A.

OpenStudy (anonymous):

heybhai ok here

OpenStudy (anonymous):

@sexybadass22 as given x intercept is 2 and y intercept is 4/5 so the slope intercept form for st. line is y2-y1=m*(x2-x1) where m is slope, (x1,y1) and (x2,y2) are co-ordinates of 2 points which is here (2,0) and (0,4/5) i.e. where the intercepts intersects axis. hence (4/5)-0=m*(0-2) m=-2/5

OpenStudy (anonymous):

sorry actual slope intercept equation is (y-y1)/(y2-y1)=(x-x1)/(x2-x1) above i wrote slope's equation only. so (y-0)/(4/5-0)=(x-2)/(0-2) or,(y)/(4/5)=(x-2)/(-2) or,(-2y)=(x-2)*(4/5) or,(-5y)=(x)/2-1 or, x+10y-2=0 is the required equation. @sexybadass22